OpenStudy (anonymous):

Perform the following multiplication. The answer should be in simplified form for radical expressions.

OpenStudy (anonymous):

\[(4+\sqrt{7})^2\] @charliealbert

OpenStudy (anonymous):

I will do it, but I want you trying to do it. Remember this: \[(a+b)^{2}=a^{2}+2ab+b^{2}\]. And tell me what you got

OpenStudy (anonymous):

\[16+8\sqrt{7}+7\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

16+8+7?

OpenStudy (anonymous):

Or: \[23+8\sqrt{7}\]

OpenStudy (anonymous):

oh, so just add the 16 & 7, got it! thanks!

OpenStudy (anonymous):

You are welcome

OpenStudy (anonymous):

Next one?

OpenStudy (anonymous):

Sorry, but I'm a little bit sleepy and tomorrow I have a presentation.

OpenStudy (anonymous):

Oh, ok...well thank you so much, goodnight ! :)

OpenStudy (anonymous):

You too